OpenStudy (imtant):

How to show that y=C1e^x+C2x is a solution for y"(1-x)+y'x-y=0?

OpenStudy (john_es):

I think the direct way: calculate y'' and y', then plug in into the differential equation. When you sum all terms the result should be zero.

OpenStudy (imtant):

Yes,but do you know how to differentiate the y?

OpenStudy (john_es):

For the first derivative (C1 and C2 are constants), \[y'=(C_1e^x+C_2x)'=C_1e^x+C_2\]I think you can follow from this point.

OpenStudy (john_es):

Do you understand how to finish the problem?

OpenStudy (imtant):

No.Can you show how you factor the x?

OpenStudy (john_es):

What do you mean by "factor the x"?

OpenStudy (imtant):

I don't understand how you get C1e^x+C2.

OpenStudy (john_es):

Ah, ok. It's all in the derivatives, \[y'=(C_1e^2+C_2x)'=(C_1e^x)'+(C_2x)'=C_1e^x+C_2\] Remembering the derivative of x is 1.

OpenStudy (john_es):

The second derivative, \[y''=C_1e^x\] And then you can plug in into the original differential equation, \[(C_1e^x)(1-x)+(C_1e^x+C_2)x-(C_1e^x+C_2x)\]

OpenStudy (imtant):

Okay.So for the second differentiate there's no need to write C2? right?

OpenStudy (john_es):

right, because (C2)'=0.

OpenStudy (imtant):

Thanks John_ES. :)

OpenStudy (john_es):

You're welcome.